OpenStudy (sleepyjess):

Determine whether the sequence converges or diverges. If it converges, give the limit. 60, 10, 5/3, 5/18, ...

OpenStudy (sleepyjess):

@perl

OpenStudy (perl):

to go from first term to second term, you divide first term by 6. to go from second term to third term, you divide second term by 6. , and the pattern continues

OpenStudy (perl):

so lets make a rule for the n'th term

OpenStudy (sleepyjess):

60*1/6\(^{n-1}\)

OpenStudy (sleepyjess):

What is the difference between a converging and diverging series? That is where I am really getting stuck, and the lessons/textbook says nothing about it.

OpenStudy (perl):

this is a sequence, not a series . a series is a sum, a sequence is a list

OpenStudy (sleepyjess):

ah, I keep getting them confused *facedesk*

OpenStudy (perl):

a sequence converges if it gets closer and closer to a specific real number. so for example this sequence converges to 2: 1.9, 1.99, 1.999, 1.9999, 1.9999, ...

OpenStudy (perl):

ok does this sequence get closer and closer to a number. lets write the decimal equivalent of the sequence? 60, 10, 5/3, 5/18, ... = 60 , 10, 3.33, 1.11, 0.37, 0.123, 0.041

OpenStudy (sleepyjess):

It is getting closer and closer to 0

OpenStudy (perl):

yes :)

OpenStudy (sleepyjess):

So that is all?

OpenStudy (perl):

a few more terms 60 , 10, 3.33, 1.11, 0.37, 0.123, 0.041, .0137, .00457, .0015, ..

OpenStudy (perl):

this is not a mathematical proof, but it is evidence that it converges

OpenStudy (perl):

if you want to be more rigorous, we can look at the limit of the nth term , defined by the rule or formula for the sequence

OpenStudy (sleepyjess):

O_O that sounds complicated

OpenStudy (perl):

you found a(n) = 60 / 6^(n-1), where n=1,2,3...

OpenStudy (perl):

what happens when n gets very large (goes to infinity) you have 60 / 6^(very large integer) = 60 / very large ~ .000000... almost zero

OpenStudy (sleepyjess):

ooohhh... that is less complicated than I thought :)

OpenStudy (perl):

or you can even use infinity notation 60 / 6^(oo-1) = 60 / (oo) ~ .000000... nearly zero

OpenStudy (perl):

this is better argument to show it converges. you can even be more rigorous using an epsilon delta proof, which i will spare the details (i think its overkill here)

OpenStudy (sleepyjess):

Thanks :D Could you help on one other?

OpenStudy (perl):

ok